I reported in the replied-to entry that the Q of the Taiwanese cantilever at low temperatures actually appeared to have gotten lower at low temperatures, relative to the case before the Si spacer was added to the clamp to avoid skewness. However, the data from the longer run over this past weekend (see the ~20-hr stretch below) seem to suggest a Q not significantly different from that measured in CRYO:1213.
Interestingly, the online phi measurement starts out at the higher level I indicated in the previous post, but then slowly approaches a level not inconsistent with the ~1.5 x 10-6 number from before the spacer addition. The title is misleading, as the system actually approached a minimum temperature of ~90 K on Saturday, but the thermoelastic noise prediction is roughly flat over this temperature band, so that shouldn't be a factor, and the associated deflection from this temperature shift should not be enough to account for this drift via calibration error.
As I discuss in the quote, I had hoped to make a continuous phi measurement as the system warmed leading up to today, but at the time I neglected to consider the thermal deflection, which over such a large temperature swing completely rasters the beam off the QPD. In retrospect, I'm lucky that this effect didn't break the cantilever as the sensing gain was reduced from the misalignment---thankfully, the loop destabilized quickly enough that the watchdog script killed the feedback before anything happened.
So, it looks like we'll have to make this measurement the old fashioned way, point-by-point, which is why I spent time reconfiguring the temperature control today. I'm running an active measurement overnight at 120 K to see if we see a Q bump there.
As you can see, the loss is hovering around 3 x 10-6, giving a Q around 3 x 105, which is slightly but significantly lower than what we measured before I added the Si spacer to avoid skewing the clamp (CRYO:1213). I would chalk this up to the spacer actually making the clamp worse, but we did in fact see a huge improvement at room temperature (CRYO:1216). So, like, what the hell man?
I've left it running to collect more data over the weekend. I haven't gone over the temperature readout/control system with Nic, so I set up a simple temperature readout in the meantime so that we can have at least a coarse Q(T) measurement as it warms. To do this, I simply put a 1k resistor inline with the RTD and put 5V across with a lab supply. The second set of RTD leads goes to the temperature readout input in the digital system, so this is now just a DC readout of the voltage across the RTD. The lockin input channel X1:SCQ-TEMPERATURE_LOCKIN_DEMOD_SIG_OUT is calibrated to volts, and is equal to 5 V * RRTD/(1k + RRTD).
It was a little unclear to me how the digital temperature control was supposed to work as it was built, so I made some modifications today.
The previous implementation used a digital lockin setup (as does the new one), but the output of this was converted to an error signal for the temperature control loop using a relatively primitive calibration, so I added some math into the model to make it a little more exact. The changes can be divided into two sections: 1) the demod voltage to RTD resistance section and 2) the RTD resistance to temperature section.
To get a nice linear temperature signal using the 4-lead sensing method, the RTD current should not be determined solely by the RTD (otherwise, the readout just sees the excitation directly). So, I have added a 1-kOhm resistor in series with the RTD in the LO path, just as I did with the temporary setup described in CRYO:1217. The series resistor and the RTD now form a voltage divider where the RTD resistance can be inferred to high precision in the limit RRTD << Ri (= 1 kOhm). This is almost always true, but the model does include the exact expression for RRTD(Vout). To set the calibration, one must enter 2 values:
For this, I have implemented the full, quartic formula from the ASTM standard (see our RTD manufacturer's page). Other than 4 preprogrammed empirical constants (and the Kelvin conversion of +273.15 at the end), this only requires one input from the user:
The heater actuation section was in pretty good shape, so I didn't really have to make any modifications there. One thing I did do was add a heater power calculation, which requires the user to enter the heater resistance.
On the hardware end, since I'm using the bigger steel block clamp which also has the higher-resistance (100-Ohm) power resistor, I found that I needed more juice than what the DAC -> voltage amp/buffer that Marie and Nic used could provide (this circuit was regulated to 15 V, giving a max power of 2.25 W, which just isn't enough). I stole my Sorensen HV supply back from the CTN lab for now, as it seems to be unused, and used it as a HV amplifier via the external control feature. Since this unit doesn't allow voltage range limiting in remote mode, I added a 1/10 divider between the DAC and it so that I didn't have to trust software limiters. Really, I should attenuate the HV output, but I couldn't think of an easy way to do that with the stuff I had on hand. Anyway, the railed heater voltage from a +10 V DAC signal is ~40 V --> 16 W.
Finally, I edited the SCQ master screen to reflect all these changes. Here, you can see the system being held at 120 K:
The main thing I wanted to do with the coherence+ beat @ 300K beat data is subtraction.
Something sort of interesting (but not unexpected when you think about it) happens when you start to combine multiple coherences in this way:
If the (quadrature) sum of errors in your estimates of coherence (mag^2(rmsav(CSD))/(PDS1*PSD2) in this case) is sizable compared to the residual noise, equation 7.33 can give you negative values of power (which corresponds to imaginary values of ASD).
I pondered for a bit and came to it is very unsurprising that subtracting two things of similar value (1, sum of coherences) to get to a smaller thing (tiny residual noise) comes with its own problems.
Here is the subtraction residual:
Multiply the calibrated beat signal by this to get the answer "what is our best upper bound on the residual noise in the beat not due to unsuppressed frequency noise?"
The 1kHz to 10kHz slope is where the PDH loops shift from being noise limited to gain limited.
The SR785 estimates coherence as follows:
We can use the coherence measurements to tell us what PDH_E/PDH_W is, which can help illuminate the earlier "what's up with the PDH calibration" problem.
sqrt(Coh(PDH_E,PLL)/Coh(PDH_W,PLL)) = ASD(PDH_E)/ASD(PDH_W)
What's going on?
PDH_E[Hz] / PDH_W[Hz] * 0.88 = PDH_E[V] / PDH_W[V], so our PDH calibrations have to reflect this ratio:
EastCal [Hz/V] / WestCal [Hz/V] = 0.88
The ratio of the calibrations I used in elog:1219 was incorrect by a factor of 2.1.
Restricting the ratio of Ecal/Wcal to be 0.88, and tuning the W PDH cal up to 5e5 Hz/V, I get:
The DC calibrations quoted in elot:1219 are incorrect for unknown reasons (My money is on "oops I forgot that I added an attenuator" error which makes the demod gain I used (0.63) incorrect.)
Transmissed RIN spectra
Trans noise level was minimized by tuning the optical PDH offset (generated via RFAM, tuned via waveplates by the EOM), and the electronic offset (tuned via low noise offset adjust knob on LB1005 box)
Measured with PDA20CS on pickoffs from each path. Trans DC voltage ~1V, TransZ gain 10 or 20 dB (not PDA20CS dark noise limited)
PDH error signal measurement:
Measured using the "error monitor" output of the LB1005 boxes
High frequency portion is multiplied by 2 b/c 50 ohm impedance of 4395 + LB err mon
OLTFs taken by measuring between two points on the loop, injecting on either side of the points, then dividing out those measurements - the extra measurement saves on calibration time sunk.
Calibrated PDH err plot vs PLL:
Something is clearly wrong with the calibrations I used for the PDH error signals. I suspect the PDH East calibration based on morphology (when I increase the E PDH calibration by ~5x, the quadrature sum of E and W error signal residuals line up with the calibrated beat signal, which is consistent with the sum of coherences being near unity (coherence measurements between PLL, PDH_W, and PDH_E indicate that unsupressed laser frequency noise is 10x higher than any other noise source at frequencies between 10kHz and 100kHz [no data on coherence above 100kHz])
The PDH calibrations can be verified/overthrown by using the measured PDH OLTFs and dividing out the known parts of the plant.
Beat raw ASD from 300K Dec measurement set:
The PLL setup was a Marconi and an SR560 after the demod/lowpass.
PLL noise was determined by replacing the transmitted beat with another marconi at the same RF power / frequency, with the VCO functionality disabled (the noise goes WAY down when you do this). Input noise of marconi we were using as a VCO was the dominant effect at high range level. Range was set as low as seemed reasonable to trade off between averaging and noise. The bucket noise is definitely not noise from the PLL readout.
OLTF was measured as G/(1-G) in each state
Control signal calibration is V_ctrl*(marconirange/1.41V)*(1-G)/G
Error signal calibration is V_err/(7e-2 V/rad)*(1-G)/1.41
Calibrated error and control signals for the beat are:
Data/plots/matlab code for processing all on svn in CryoLab/Measurements/CoherentSub/ directory
The cantilever had cooled to around 100 K by this morning, so I set up the mode ringer and began an active measurement on the fundamental mode. The online loss angle measurement for a 3-hr period beginning around an hour after lock is shown below (this is the control signal filtered by a 2nd-order low pass at 0.2 mHz.
The LN2 dewar was refilled today, so I filled the cryostat and we'll see how it looks at low temperature tomorrow.
On Tuesday night, when I added the Si spacer to the clamp, I measured a Q of ~46000, but I noted that it had been increasing up to that point, likely due to the residual gas damping (see CRYO:1214). Last night, I made another measurement and found it to be much higher, at ~1.2 x 105 (tau ~ 350 s). This is much better than we have seen at room temperature thus far, so it looks like my spacer addition has helped.
I remeasured this an hour or so later and saw no appreciable increase. I checked again today and it appears as though it may have increased slightly, but it was hard to say for sure due to higher environmental noise. Really, we need the steady-state ringdown to make a good measurement at this level.
I called the campus service yesterday morning to have the LN2 dewar refilled. They got around to it today. New dewar number is 102.
The most recent measurements on the Taiwan-sourced Glasgow-style cantilver (see CRYO:1213) are encouraging, but the best Q measurement at low temperature is still a couple orders of magnitude worse than what is theoretically achievable, and about one order of magnitude worse than our conservative clamp loss estimates. Also, I've done some measurements on other modes (that have different expected clamp loss contributions due to the relative strain energy ratios) to try and sort out what is going on, with little success. Finally, some modes---including the 2nd bending mode at ~650 Hz---exhibited very low Q for no known reason.
One thing I thought about is that, since the Taiwan cantilever did not fit in the groove that was built into the block for the Glasgow-style cantilevers and therefore is just sandwiched between the two large pieces making up the clamp (see CRYO:1211), the clamp is likely pushing down at somewhat of an angle, which could lead to all sorts of non-idealities. Since the other Si samples we have lying around are roughly the size of the clamping region of this cantilever (~300-500 um), I opened up the cryostat today and reclamped the cantilever using a spare broken-off 300-um-thick cantilever piece as a spacer on the other side:
Pumping it all back down, I immediately measured Qs a bit higher than what we saw last time around at room temperature. The last measurement I made before leaving was tau ~ 135 s ==> Q ~ 46000, though it had been increasing up to that point, likely from the residual pressure, which was at ~10-3 Torr when I left. Compare this with the Q of ~14000 from the last time around, though admittedly I did not record the pressure at which this was measured.
We measured the Q of the fundamental (~106 Hz) mode of the Taiwan cantilever in two ways. First, we used Nic's active steady-state method, and then we did a traditional ringdown. The results seem to agree, but the precision of the first method is much better due to the dynamic range of the readout for this mode: the motion becomes nonlinear at an amplitude only a few times greater than the background excitation level. Over a ~4-hr average, the loss is measured to be 1.45 x 10-6 ± 2.9 x 10-7, giving a Q of ~6.9 x 105.
Here is a plot of the instantaneous phi from the calibrated control signal. This data has already been fed through a ~1-hr lowpass, and then the data from the initial settling time has been truncated away. The mean and standard deviation of the rest of the points are what is reported.
After this measurement was made, we shut off the servo and allowed the mode to ring down. Here is that ringdown, along with a predicted range of theoretical curves using the result from above. As you can see, they are fairly consistent with what is measured, considering that the system quickly reaches a regime where it is excited by the environment (that is, only the initial part of the ringdown, where the agreement is good, is very trustworthy).
This Q is a couple orders of magnitude lower than what is expected for this mode at this temperature, but it is also only a factor of 2-3 worse than the best measurements using a similar apparatus at Glasgow (to my knowledge).
It bugs me that we don't seem to have any information about what steel looks like at low temperatures. Given my COMSOL strain energy modeling, the energy ratio for this mode is about 3 x 10-4, so this could be explained by clamp loss if the steel Q is as low as a few hundred. I'm looking into other modes to try and support or refute this hypothesis; since different modes have different energy ratios, we may be able to see what's going on. In parallel, I'm asking Matt and others to find out what is really known about cryogenic steel.
Given that we see some excess noise in our 3-mm Laser Components diodes (IG17X3000G1i), especially with one of them, I went ahead and did a more careful measurement of both the dark currents and noise.
To make this measurement, I switched to OPA140 transimpedance amps on the M2 readout board and used a 1-M transimpedance.
The OPA140 has a bias/offset current of 10 pA max and an offset voltage of 120 uV max, the latter which therefore limits the DC current sensitivity with this transimpedance to 0.12 nA. There is also some allowed current variation over temperature (±3 nA bias and ±1 nA offset over -40 to +125 °C), so this can add some more DC uncertainty if the lab temperature is a few degrees away from 25 °C. Plugging the outputs into the DMM without the diodes connected, I measured 0.0 mV and 0.2 mV on amps 1 and 2, respectively. This is consistent with the op amp spec.
The Laser Components spec for the dark current (at 5-V bias, where I measured it) is 20 nA typ, 100 nA max. Plugging in the diodes while keeping them in a blocked box and with the room lights off, I measured the following bias currents (output voltage divided by 1 M):
The first 3 diodes are within the max spec, while 7845 seems to be exhibiting some catastrophic failure mode where the dark current is avalanching whenever the bias is engaged. Below is a plot of the measured amplifier output after a turn-on of this diode, with one of a healthy diode for comparison to the right. This was taken in the middle of the testing, and the last measurement of the current before this turn-on was around 140 nA. As you can see, there is an initial slew (not inconsistent with the timescale of the bias turn-on), followed by a slow but monotonic increase of the dark current over time. When this was repeated, the initial slew brought the current again to the last-known highest level.
I used the same transimpedance amp setup to measure the noise. All diodes show spectactularly higher noise than the advertised level of 3.2 x 10-14 W/rtHz NEP (~3 x 10-14 A/rtHz), with a 1/f characteristic that, if extrapolated, would not intercept the quoted spec until ~1 MHz. A frequency for this spec is not mentioned on the datasheet. In all cases, the circuit was allowed to equilibrate for a few minutes before a measurement was made. The spectra below were found to be stationary, with the exception of occasional glitches.
The readout noise is limited from several 100 mHz up to near 1 kHz by the Johnson noise of the 1-M transimpedance resistor, above which there is some noise peaking centered around 40 kHz that is not inconsistent with other measurements I have made with this op amp in very-high-impedance environments (c.f., 40m:8151).
IF the DC dark current is out of spec, we might be able to get a replacement. Might be specs on the website. I think Frank had a Keithley instrument to measure dark currents that are low - probably in his diode destruction elogs or DCC docs.
Nic got a Glasgow-style cantilever from a group in Taiwan, and a quick test in the rapid cycle chamber showed that it had pretty low loss, so we are running it in the cryostat now. As a reminder, these are the rough dimensions of this style cantilever:
Below is a photo of the box it came in, showing the actual 92-um thickness of this sample, as well as a shot of it in the vacuum chamber. For some reason, this particular sample's clamping tab did not fit in the groove that Nic had built into the clamping block for the other Glasgow cantilevers, so I had to mount it to the side against the flat faces of the clamp (as I've been doing with our larger samples).
This evening, we transferred it over to the cryostat and restored all the electrical connections for what will hopefully be a fruitful cryo run. Here is a ringdown of the fundamental mode (~106 Hz) at room temperature:
The measured decay time of 41 seconds corresponds to a Q of around 14,000, which is about as good as we expect at room temperature. This sample is probably better than our other ones for at least 2 reasons:
Given that we didn't see much improvement at all with our other samples when going to low temperature, I believe (2) is by far the biggest effect. The Glasgow wafers only have the clamp-region thickness extended to one side, which is modelled to be worse than if you go both ways, but it is still much better than we can do with our discrete sandwiching.
I filled the LN2 reservoir and the volume is cooling overnight. I did some rough ringdowns at a point when the steel block was registering around 160 K and found greatly improved Qs already (approaching and perhaps exceeding 105). We will continue to make measurements tomorrow.
Tonight I succeeded in using the M2 ISS readout board and the 3-mm diodes to do some real intensity stabilization using the SiFi test setup.
First, I built a foam box to use as a temporary enclosure for the PMC and diodes until we get our real box finished:
There are holes for the input and REFL beams, and the diodes are held with makeshift mounts that clamp down on the sockets. Clearly, these aren't as stiff or stable as what we're having built, but they do the job for now. There is a steering mirror before the 50/50 BS so that the position on each diode can be adjusted separately with ease.
I didn't want to open the Chachi ISS box can of worms yet, so I just built my own temporary breadboard circuit. I had done some preliminary SR560 locking, so I knew roughly what I wanted, and I measured the modulator -> PD transfer function again today and verified that it was flat well above 100 kHz. I made a 2-stage pole/zero-style circuit, with a double (removable) pole at 300 Hz and a zero at 10 kHz to bring phase back to -90° around the target UGF of 100 kHz. It looks like this:
I wanted to DC couple, so I came up with an idea to pick off the stable 5-V bias supply from the M2 board and sum it with the (negative) output of the in-loop PD in the first stage of the servo. I had some current-related issues with the summer at first, but these went away when I increased the input resistors a bit (n.b., to fix the gain I had to change some other components, and as a result the controller TF is actually slightly different than shown above, but not much).
Hooking it up, it locked right away with the expected UGF of near 100 kHz (not yet measured, but inferred from the transfer functions and spectra). Here is the stabilization result:
As you can see, the out-of-loop signal is stabilized to the shot noise level (which is higher than the bare shot noise for half the beam due to the well-understood correlated noise imprinted by the loop) from about 5 kHz down to just below 100 Hz. Below this, there is clearly some differential environmental noise between the PDs. I did some beam scanning to try and minimize with some success, but not much. I'm not sure what the coherence below 20 Hz indicates---the in-loop signal is suppressed to below the measurement noise level, while the OOL signal exhibits excess differential noise, so I don't see why there should be any coherence.
In any case, this is a nice verification that:
Frank had = PeterK had = It went to LHO
I wonder if it helps to use FEMTO's DLPCA-200 that I have somewhere in my lab.
I was preparing to do an initial test of the M2 ISS readout board with the 3-mm diodes on the SiFi test setup when I noticed some anomalously high noise on one of the diodes. So, I decided to make a more careful measurement and test all 4 diodes. I found that only one (S/N 7845) exhibits this very bad excess 1/f noise, but all four have it present at some level.
For this test, I had the transimpedance fairly high at Z = 2.7 k since I am only working with < 5 mW of power, and the diodes were completely blocked for this measurement and put in a dark box. The bias was 10 V at first, but then reduced to 5 V in an attempt to reduce the excess noise after I read on the datasheet that 10 V was an absolute maximum for some reason. I did not record the difference in noise from 10 V to 5 V, but this is a test I will likely try (though perhaps not up to 10 V anymore).
While 7845 is clearly bad, the others are probably OK for now; they are not acceptable for low-power/high-Z operation, but are likely just fine for our high-power testing since we will expect shot noise levels of >100 pA/rtHz, with SNR with respect to PD noise increasing as .
I wanted to do some intensity feedback testing, for two reasons:
So that I can have as much power as possible, I removed the fiber phase modulator and installed the amplitude modulator in its place. To generate the PDH sidebands, I simply drove into the laser bias tee with the 30 MHz oscillator signal and increased the amplitude until I got the same modulation depth as I measured with the modulator. I also had to readjust the demod phase via cable lengths, but after that the cavity locked just as before (and with an identical OLTF---not shown here). I don't claim that this locking technique is as good as using a phase modulator, in light of possible RFAM effects, but it is likely fine for intensity testing.
I also tried to increase the DC drive current of the laser, but it kept stalling after I tried to increase it above ~115 mA (the output power would increase in accordance with the plot on the datasheet, but then would suddenly crash and not return if the current was lowered until the driver output ON/OFF was cycled---not sure what gives here). So, I set it to 100 mA, where it seemed stable. The output of the laser head at this current is ~12 mW, so the max-transmission output of the amplitude modulator is about 6 mW (due to the 50% insertion loss). Adding a slight DC offset to the modulator, I reduced the output to ~92% to get some linear actuation strength for feedback.
I then tried to create an AC-coupled loop with an SR560, but had problems with stability on the low end. Eventually, I gave up and used the A-B function to subtract the measured DC level of around 4 V from the TRANS PD signal. I then put a pole at 300 Hz and scaled up the gain until I saw oscillations up near 100 kHz, and then slightly back down. Using this offset-subtracted DC-coupled loop, I was able to get solid in-loop performance, obtaining a UGF near 100 kHz and suppressing fluctuations to the dark noise level (consistent with the PDA255's noise) over a wide band.
The next step will be to use my low-noise readout optoelectronics and try out the Chachi servo.
Using the REFL PDH setup I built the other day (and that was detailed somewhat by Nic and Chris W. in CRYO:1204), I calibrated the error response so that I could make some further measurements. To refresh, this is using 30-MHz sidebands applied using one of our fiber phase modulators, sensing with a 1611 in reflection. The sideband drive was 0 dBm.
Using the sidebands as a reference, I calculated the slope at 6.4 nV/Hz:
Note that the error signal is slightly asymmetric, but there is no large offset.
With this information, I made some measurements:
I wanted to measure the actuation transfer functions afforded by:
All measurements were done with the west laser head, driving as described and reading out at the error point, then correcting for the loop gain and calibrating to Hz.
Everything is more or less as expected:
I now have another calibrated measure of laser frequency noise, wherever it dominates over the PMC length noise. I measured the error signal, corrected for the loop gain and calibrated to Hz. For comparison, I've added the measurement using the Zurich PLL on the beat between the two free-running lasers on 12/17/2014 (see CRYO:1185), as well as the RIO spec for this laser.
As you can see, tonight's measurement agrees quite well with the earlier one upt to ~1 kHz, above which the old measurement is probably marred by the relatively low-bandwidth PLL. It seems that the PMC is quiet enough to see the laser noise throughout, and the new measurement now sits closer to the spec up to the highest available point at 10 kHz. Below ~50 Hz, we are probably seeing the well-documented excess noise from the ThorLabs driver. Everything looks as expected.
Relatively early in the night, after having measured the actuation transfer functions, I sucessfully locked the cavity via feedback to the laser (as opposed to the PMC PZT) for the first time. Below is a comparison of the OLTFs for a 1-kHz loop using the same servo shape (a pole at 1 Hz) using both actuation schemes.
Because of what I had hooked up at the time, I only did this with the (low-bandwidth) LDC201C, so, while the absence of a ~10-kHz resonance is clear, the phase margin is not improved at all (worsened, actually). I only report this as a milestone, and the margin afforded by the ITC502 or by directly driving via the bias tee should be far better.
In order to better measure the effect of this nonlinear current to frequency modulation, we'll need to do Zach's measurement but with much higher drive frequencies. (His measurement was 1kHz).
We'd like to do a full TF of the nonlinear current amplitude modulation path to the laser frequency. There are two effects in Zach's setup that limit the bandwidth of the measurement.
First, is the modulation input of the Marconi, which only reaches 30kHz. We plan to use a mixer to do higher frequency AM of the RF carrier.
The second is the frequency readout. We potentially could PLL the two lasers together and have a pretty high bandwidth readout. or, instead we decided to add some additional PDH sidebands to the light using the fiber modulator. This was then sensed in reflection of the PMC and demodulated. We used 30MHz at 0.5Vpp into the fiber modulator.
With this setup, we were able to measure some amount of nonlinear current to frequency modulation, and when we unlocked the cavity the transfer function was reduced by at least 20dB, which rules out some other coupling path.
Next step is to set up high bandwidth AM of the 500MHz marconi output (driving the current).
Nic elucidated to me today Chris W.'s idea for getting truly wideband (~500 MHz) actuation out of our diode lasers. In case the reader isn't familiar, the lasers have two parallel linear actuation pathways converting current into frequency: one from current modulating the temperature, which is the strongest effect at DC and then dies off above ~1 MHz due most likely to the thermal response, and another, weaker but much wider-band, flat pathway arising from solid state effects that did not survive the elucidating. At some frequency (around 50 MHz, I believe?), there is a crossover between these paths, but there is a differing sign, which creates a "non-minimal-phase zero", leaving the phase at -180° and making the overall system a difficult actuator to deal with at high frequencies.
As I understand it, Chris's idea involves using the full, nonlinear current-to-temperature response to effectively circumvent the direct linear response at low frequencies. This can be done, for example, by pumping a strong RF carrier current (say, around 1 GHz) into the diode, and then using amplitude modulation on this carrier to produce baseband frequency actuation from the temperature beating. By choosing the phase of the AM correctly, one can make it so this pathway (now dominant at low frequencies) results in a nicer crossover with linear pathway #2 from above.
I performed a very simple proof-of-principle test today by doing the following:
Trimming the RF amplitude and phase a bit to get a nice result, I was able to take the two spectra shown below. In the first trace, only the direct current line is present at 1 kHz. In the second one, the RF source is engaged and you can see an exact cancellation of the line in the error signal. Increasing or decreasing the RF (or audio) amplitudes led to the reemergence of the line (assuredly with 180º relative phase from one case to the other). To do the wideband actuation, one would simply make sure that the RF power is strong enough that the nonlinear path dominates.
Following my preliminary conclusion from yesterday (CRYO:1201), I set out to confirm or deny this seeming decrease in Q for a given clamp when going from the simple vacuum chamber to the cryostat.
One potential source of extra damping I considered was the wires attached to the block for the power resistor and RTD, so, while I still had the clamp in the cryostat assembly, I just disconnected these wires and pumped down the cryostat to see if I saw an improvement. I did see an increase in Q from ~3000 to ~5500, but not to the full 7000 I saw before in the standalone chamber. So, I conclude that there is some appreciable damping added by this kapton wiring. We need to use less rigid wire for the last stretch between the coldplate-mounted strain releif and the block.
The last step was to transport the clamp back into the simple chamber and see if I could recover the Q of 7000 that I measured initially. I did, completing the circle of repeatablility. I'm not sure what else could be causing the excess damping in the cryostat.
It is a shame, because I would be very interested to see what this particular silicon sandwich clamp looks like at 120 K, but I seem to have now way of doing so without the extra losses empirically associated with putting it in the cryostat.
A lot of things happened tonight (mostly in the realm of setbacks followed by recovering frome them), but the take-home is that the measured Q of my silicon sandwich clamp seems consistently lower when measured in the cryostat, compared to in the new chamber from the gyro. Here's a rundown of what happened today/tonight:
So, it seems that the Q is repeatably lower for a particular clamp in the cryostat vs. in the simple chamber. To be sure, I'm going to do the final step of returning the clamp back to the simple chamber tomorrow and see if I again get a higher Q.
I'm not exactly sure why this could be happening. The only mechanical differences from one chamber to the other are:
I'm tempted to think that (2) could be causing some excess damping, so one thing I will try is simply not connecting these just to see if that makes the probem go away.
As I planned yesterday (CRYO:1199), I tried out a new clamp using spare pieces of broken silicon instead of sapphire washers to sandwich the cantilever (as with the last run, I used the old, stiff rectangular block clamp---the newer cylindrical one is still in the cryostat).
I didn't take a photo, but this was basically just a sandwich consisting of the cantilever (still attached to the central wafer region) as the meat and two scrap broken-off cantilevers on each side as the bread. This was all put near the center of the steel block clamp so that the clamping force was normal, and I made sure that the protruding cantilever had enough room not to be clipped by the block as it swings.
I put it in the new chamber and pumped down, and immediately measured a fairly high Q of ~6800 (ringdown tau ~ 6.4 s, while the mode frequency is ~340 Hz---slightly higher than before due to the clamping being a bit further along the cantilever).
This is the highest room-temperature Q I've yet measured, beating the ~4300 I measured after we first installed the sapphire washers on the newer cylindrical clamp (see CRYO:1191), and is within a factor of 2 of Marie's prediction in the absence of clamping loss (also shown in that post). This is also by far the cleanest ringdown I've seen: there are a few high-frequency modes present when I first deliver the impulse, but they die away and do not return. The Q also seems far less amplitude-dependent than I've noticed before.
It is a little tedious waiting for a full cryo cycle to iterate on the clamp. Also, in many cases we can learn a lot from just running at room temperature, but opening and closing the cryostat to get at the experiment takes a fair bit of effort. So, tonight I repurposed one of the gyro corner chambers to serve as a rapid-iteration room-temperature testbed. I used the northeast chamber since it had the pump connection. It has 2 KF flanges (on which I have put blanks) and 2 CF (one which goes to the gauges and valve, and the other which used to have a blank that I have replaced with a window).
I set it up next to the cryostat so that we only have to move 2 mirrors to switch between setups.
Given my revelation about the energy leakage and PEEK loss last night (see CRYO:1198), I resurrected the old rectangular block clamp to try a new idea. Namely, I just tried sandwiching the silicon cantilever (the central region with the hole, that is) between two sapphire washers, and then clamping the whole sandwich using the block clamp. The block clamp also has a PEEK base, but it should have provided a much stiffer clamp than the newer, cylindrical one, and that should result in less energy getting to the base. Here is what it looked like:
I pumped the chamber down and took a quick ringdown measurement. Unfortunately, the result was a Q in the ~2000 region, similar to what it was when we first installed the sapphire washers in the newer clamp and the bottom one was sitting on the clamp's lip (see CRYO:1191). Never fear---I have a new suspect: in looking at my photos, I'm noticing that the sapphire washers are not particularly flat. This could mean that the clamp contact is some strange shape and/or that the silicon is being stressed in some strange way.
Instead of the washers, I think I'm going to try sandwiching the cantilever between some other spare pieces of silicon that we have. If I use enough pieces to make a decently thick clamping region, this should serve the same purpose that we hoped the sapphire washers would. I'll try this tomorrow.
I sealed the cryostat vacuum line so I could use the pump for the new chamber. The LN2 reservoir was empty before I did so, and the clamp was registering around 250 K when I left. In any case, I'm going to keep iterating with the new chamber, and I think we shouldn't bother with the cryostat again until we can demonstrate a Q of close 104 at room temperature.
The cantilever was fully cooled by the time I got in this afternoon. I measured some quick ringdowns by looking at the amplitude on the scope, and estimated a Q of 2-2.5 x 104. This is slightly better than what I measured the other day before improving the clamping (see CRYO:1193), but not good---still a few orders of magnitude below what we expect. I heated the system up near 120 K and found a slight reduction in Q.
Unlike before, I noticed a strange sort of sloshing of energy into a higher-frequency mode (~1350 Hz). It was hard to tell, but I got the sense that energy was being dissipated out of the fundamental mode through this higher-order one. I looked at a time-lapse spectrum of the ringdown, and it seemed to confirm this effect. If you look at the movie below (which is just about real time), you can see that the RMS of the two modes between 1-2 kHz pump up and down, while the fundamental mode around 215 Hz monotonically decreases. If you squint, it appears that the full RMS stays constant in most cases while the high-frequency modes ring up, while they all decrease together. This, coupled with the fact that everything rings down to zero if left alone, indicates to me that energy is leaking from the fundamental mode out through these others. As an order-of-magnitude estimate, the amount of energy pumped through these modes as the amplitudes increase and decrease is not inconsistent with the energy lost from the fundamental based on the observed Q.
I did some COMSOLing to try and figure out what is going on, and at first I couldn't explain it; it appeared that even the higher-frequency modes should have too little strain energy density leakage into the steel to explain the effect, especially with the sapphire spacers. In looking a little more carefully, though, I realized that we have not been careful enough in modeling our system: at the bottom of the clamp stack, there is a PEEK platform between the clamp post and the cold plate. This is there by design, to thermally insulate the clamp from the bath (for heating), but it also considerably softens the contact there.
This PEEK piece shouldn't have much of an effect on the fundamental mode, as the energy ratio for that mode is of order 10-4. The second mode at 1350 Hz is nearly as well isolated. However, for the third mode around 1800 Hz, something like 70%(!!) of the energy is expected to reside in the PEEK layer. Since PEEK has very high loss, this is not good. Here are some COMSOL screenshots, with the first 3 showing the first 3 mode shapes, and the fourth showing the (log) strain energy density for the 3rd mode. Note that this model is run at room temperature, so the eigenfrequencies are somewhat higher than in my spectra.
I vented the chamber today to redo the clamping and investigate our wiring issues.
Since I observed relatively low Q even at cryogenic temperatures, I assumed there was some jankiness with how we clamped the cantilever when we installed the sapphire washers. Recall that the lower part of the steel clamp has a circular lip near the center around the screw hole, and it was too wide to allow the sapphire washers to fit around it. This meant that the lower washer was only being held by this lip, and not by the full surface area of the clamp. Also, when we installed the washers, we didn't remove the smallest can around the physics package, so we were doing a bit of guesswork as to how well aligned the entire clamp stack was. This meant that there could have been some slight rubbing, for example. Here is a photo of what it looked like in profile when I did remove the can today:
You can see what I mean about the lip, and it's also clear that the stack was not very well aligned. To fix the lip problem, I found a steel washer that was just about the right thickness and drilled the center hole out wide enough that it fit around the lip. This way, the lower sapphire washer will be supported by a larger surface from below (of course, the real solution will be to either design a new clamp or get wide-enough-ID sapphire washers). The picture on the left below shows the washer around the lip.
There was also some dust and other gunk visible to the eye, so I thoroughly cleaned all parts in the stack with methanol and isopropanol. I then carefully restacked the components and reclamped (a little tighter than we did last time, as well). The final stack is shown below at right.
I checked each connection from the feedthrough to the heater or RTD, and found that everything seemed to be in order, so there must have just been a short when we closed up last time. I wrapped some extra kapton around each connector solder joint to provide insulation and extra strain relief, and everything stayed as it should be when I resealed the chamber. I *did* accidentally break a joint on the wire for the ESD while closing up---whoops---but I decided it was more hassle to fix it than necessary for this next run. I'll resolder it when we cycle again.
The chamber is under vacuum now and I filled the reservoir with nitrogen. The clamp was at 200 K when I left around 10pm, so I'm hoping things will be calm and cool when I come in tomorrow.
After rebuilding the PMC setup (see CRYO:1195), I was finally able to move on to characterizing the Photline fiber-coupled phase modulators we will be using (MPX-LN-0.1 --- datasheet attached nope google it yourself). I measured a couple things:
As with the amplitude modulators (see CRYO:1187), I determined this simply by measusing the power straight out of the laser, then quickly connecting each phase modulator (one at a time) between the laser and the output coupler and measuring again. As I mentioned in the linked post, this is not an exact science due to the somewhat unpredictable behavior from connector to connector. Nevertheless, one can be confident at the one-to-few-percent level.
2.66 mW out / 5.00 mW in --> loss ~ 2.74 dB
2.88 mW out / 5.38 mW in --> loss ~ 2.71 dB
Supposedly, we had these two units hand selected for loss < 2.5 dB (for free, after we paid for the $500 low-loss selection of the amplitude modulators), while the standard typical loss from the datasheet is closer to what we have at 2.7 dB. An extra 0.2 dB isn't going to break the bank, but it's a bit disappointing that they didn't give us what they said. Probably too late to say anything anyway...
My plan was to use the modulators to pump light into RF sidebands, then use the frequency selectivity of the PMC to measure the SB power and back out the actuation strength (Vpi). I was able to do this, to a degree, but I was thwarted by an unexpected issue: the modulators and the fibers coupling to/from them appear to change the output mode emerging from the collimator. What's worse, the mode seems highly sensitive to any touching of the fiber whatsoever. This was most egregious with S/N 10, with which my new cavity coupling maxed out at 83%(!), even after slight empirical MMT tweaking. S/N 2 wasn't as nasty; I got ~91.5% with it.
Given this, my new plan was to make a quick-and-dirty measurement in the following way:
Measured amplitude to double REFL power: 0.78 Vpp --> 0.39 Vpk.
2*J1^2 = 17% --> gamma = 0.611
Vpi = 0.39 * (pi / 0.611) ~ 2.00 V
Measured amplitude to double REFL power: 0.52 Vpp --> 0.26 Vpk.
2*J1^2 = 8.5% --> gamma = 0.422
Vpi = 0.26 * (pi / 0.422) ~ 1.93 V
The datasheet claims 3.5 V typical, so this seems pretty good (though the spec is only officially at 50 kHz drive). Holding the amplitudes constant, I also swept the frequency down from 30 MHz to 10 MHz, and the reflected power was stable to around 5%.
Again, this is only really a quick-and-dirty measurement. Unfortunately, the only real way to get a good measurement is to reprofile the beam again with each modulator in place. Then, the contrast defect can presumably be brought down closer to 2% or better again, and the measurement can be made more cleanly. I'm hesitant to waste time doing so, though, given the observed mode dependence on the fiber resting position.
I was having some issues with the beam(s) I had previously mode matched into the PMC. Apart from not having gotten great coupling to begin with, the alignment seemed to have drifted over a few days (I noticed this last week). I attributed this to 2 things: 1) the MMT I had was a pretty sensitive one, owing partly to the fact that I had to work with the beam far outside the Rayleigh zone due to the beam beat recombination being upstream, and 2) having the recombining BS in the way, I was susceptible to clipping in the output path I was using for the PMC. I don't really need the beat setup at the moment, and I can do the modulator characterization using a single laser, so I decided to rebuild the PMC test setup using a single laser.
As a first step, I simply remeasured the output beam profile of the West laser using the razor blade technique. The beam seems very circular and not astigmatic, so I only profiled in the horizontal direction. The result:
Using this, I recalculated a better MMT:
w0x = 303.7849 um
w0y = 303.7849 um
lens 1: f = 103.2118 mm
lens 2: f = 206.4236 mm
d1 = 6.161 cm
d2 = 14.3007 cm
d3 = 29.5383 cm
(Total distance = 50 cm)
I then installed this, aligned the PMC and was able to get ~96% coupling with little trouble. By locally optimizing the second lens, I pushed this to about 97.5%. While a bullseye was faintly evident on the card in the first case, it was very hard to tell what was reflected after the reoptimization.
I borrowed the RF electronics from the steel gyro PMC temporarily (splitter, mixer, bias tee and filters). For some reason, the 1-MHz dither I used with that PMC did not work with this one, but I was able to derive a nice error signal using a 300-kHz dither at 3 Vpp. I wanted to use the uPDH box I used to use before I had the digital servo for the gyro PMC, but I forgot that Eric Q had borrowed it for the 40m. Instead, I was actually able to lock robustly and stably with just an SR560 and a single pole at 10 Hz. The control signal stays within its output range over ~10 min+ time scales. (I didn't bother measuring the loop---all I needed for my phase modulator characterization is essentially a DC lock, and the bandwith was easily 10s-100s of Hz).
The transmission dither lock leaves the REFL port open so that I can measure the rejected sideband light pumped by the modulator as planned.
I monitored the reservoir level periodically over the day and night. As of the evening, there appeared to be ~1 cm of LN2 still there. As of around 4am, it appears empty, so it should be OK to open tomorrow. I've sealed the vacuum and shut off the pump in preparation.
I'm going to wait for things to warm up and then vent the chamber so that we can:
Dmass helped me solve the Great Funnel Problem of 2015 by fashioning a foil extender to put in the tip of his metal funnel, since my glass funnel has a spout that is too narrow to get enough nitrogren through it. We spent some time yesterday afternoon filling the reservoir, after which I waited and then came back to see if it was still holding liquid. It was, so I added some more and left it overnight, and there still seemed to be some liquid by late this afternoon.
Assuming the cold volume had had enough time to reach low temperature, I made a quick ringdown measurement, only to find that the Q had only increased from ~4000 to ~8000 between room temperature and now. I think this means that the clamp integrity afforded by the sapphire washer sitting on just the lip of the steel clamp is not good.
I ordered a new LN2 dewar and it has just arrived. Appropriately, for me, it is #305.
Yesterday, we opened up the small cryostat and installed the sapphire washers (SwissJewel SP-175). This is hypothesized to increase the resonator Q by reducing the strain energy leaking into the lower-Q steel clamp.
We found that the inner diameter of the washers is slightly too small to accomodate the inner lip of the lower part of the clamp. We were able to make do just by having the lower sapphire washer sitting on this lip---rather than on the full wider area of the lower clamp section---but it is not ideal.
Nevertheless, we clamped it, resealed and pumped the chamber down. As it pumped, I rebuilt the HeNe optical lever readout. When I finished, I was quickly able to tap the cryostat and see a mode ringing at almost exactly 250 Hz, which is known to be the frequency of this cantilever at room temperature. At a respectable pressure of several x 10-5 Torr, I made a quick-and-dirty ringdown measurement using a scope and a stopwatch. I estimated at roughly 2.5 seconds, giving Q ~ 2000. This was already a few times higher than Marie was able to measure at room temperature (see below).
I went down today and did an actual measurment, using the Zurich box sampling at 7 kHz as DAQ. Fitting the envelope by eye, I found a time constant closer to = 5.55 s, giving Q ~ 4300 (I don't think my stopwatch method was all that wrong yesterday, but I do think the residual gas might have been contributing at the time---the pressure is now at 10-7 Torr). This is not only much better than the previous result, but also within a factor of less than 3 of the expected result for Si, according to Marie's data. Given how cavalier we were with the clamping, I'm fairly confident that the sapphire washer idea (and therefore also the monolithic thicker-clamp idea) works as intended.
To continue with the laser/modulator testing, I have added Dmass's old PMC to the temporary characterization setup. I have used the other output of the 50/50 BS that combines the two laser diode outputs, so that we can keep the beat setup intact while also being able to send either of the two beams into the PMC.
To do this, I:
The coupling isn't stellar yet, at roughly ~66%, but the MMT is fairly tight and I'm sure I can improve easily. The laser and cavity are stable to well within a linewidth at high frequencies, and only drift apart over many seconds.
Some things I plan to do with this setup:
The boring way:
Tonight, I did some characterization of the Photline fiber-coupled amplitude modulators we will use for our experiment (MXAN-LN-10 --- datasheet attached nope google it yourself). These are electro-optic devices that work by using an internal mach-zehnder to convert phase modulation into amplitude modulation.
The test setup for all measurements was the same. I used the exact configuration that I have been using for the beat (see CRYO:1182), but I simply blocked one laser, so that only one beam was hitting the 1811 PD. The amplitude modulators were inserted (one at a time) between the East laser and its output coupler.
The first thing I did was to investigate the insertion loss of the modulators. We chose the low-loss option, which just meant that the company hand-selected modulators with loss of < 3dB (= 50% power transmission).
I didn't go crazy with precision here, because systematics with fiber coupling can easily prevent a measurement to better than a few percent (an example of this: I installed a 1-meter patch fiber between the laser and the output coupler, instead of the modulator, and I actually saw a slight increase in output power vs. the case with the laser going straight to the output coupler… go figure).
In both cases, I measured very nearly 50% reduction in power (at the top of the MZ fringe---see below) vs. the case with no modulator. So, these things have a loss very close to 3 dB, as advertised. An important thing to point out is that we will need to bias these away from maximum transmission to get a linear PM -> AM coupling, so the real power reduction in our setup will be more than 50%.
These modulators have an SMA-connectorized "RF" input, as well as two bare pins connected to a separate set of "DC" electrodes (they also have two more pins connected to the cathode and anode of an internal PD, presumably at the other MZ output port, which is kind of cool). As far as I can tell, the RF input is also DC coupled, only it is 50-ohm terminated.
I did a DC sweep of both electrodes from 0-10 V while measuring the output power:
(The RF applied voltage range is lower due to sagging from the 50-ohm load).
Fitting these curves, I determined the following Vpis:
These are consistent with the numbers listed on the datasheet.
Next I measured the actuation transfer functions ([RIN/V]) from 1 Hz to 100 MHz, driving the RF input while applying a mid-fringe bias to the DC input, and using
Note the dead zone from 50-500 kHz---this was by accident, as I forgot to check the low-frequency resolution of the RF measurement. I will redo this sometime.
Here are the results:
The response very flat, and roughly what is expected from the DC sweep:
(1/P0) * dP/dV|mid-fringe = pi/Vpi ~ 0.5 ( = -6 dB).
I upped the sample rate of the x1cry model to 64K, in the following way:
The only tricky part is the last step. Changing the sample rate requires the filter coefficients to be updated, so they still match the filter designs. But when you open the filter file in Foton, it does the opposite: updates the designs so they match the old, incorrect coefficients. Since x1cry had only a few filters defined, I went through the file and reverted the designs by hand. (Newer versions of Foton would let you automate this step.)
I spent some time tonight measuring the free-running laser beat noise in various ways. Recall that, as of yesterday, I had tried setting up a couple analog PLLs to no avail and I didn't trust the spectrum I was getting from the Zurich PLL. So, I wanted to measure it another way to see if I could corroborate.
First, eye candy:
Now, an explanation of the various measurements.
I-Q demodulation method
This is a method I have used with some success in measuring the Marconi noise in its quietest state (with no modulation and therefore no means of feedback---see ATF:1877). It is done in the following way:
The main complication here is that, as you can see in the plot, the high-frequency RMS of the beat is several tens of kHz, which means you still have to sample at a high rate to get what you need. The best acquisition scheme I could think of was the Zurich box, which can do 460 kS/s. Still, to take meaningful data, I had to very carefully tune the laser beat to the Marconi LO and then quickly engage acquisition before the (wildly fluctuating) IF signals drifted above the Nyquist frequency (around one second of data was used to make this trace).
That said, the result doesn't look that crazy, and in fact it agrees very well with the DFD measurement that was carried out in a completely different way (see below).
Delay-line frequency discriminator (DFD) method
This is the usual scheme where one mixes a signal with a time-delayed version of itself to create dispersion. What I did:
This method worked swimmingly and reproduced exactly the result I found using the I-Q scheme. The noise floor (cyan in the plot) was measured by sending a quiet Marconi sine wave of the same amplitude and frequency as the beat through the pipeline.
Zurich PLL method
This method is incredibly straightforward. Simply plug the beat (ensuring it's < 1 Vrms and under 50 MHz) into the Zurich box and lock the internal PLL by pressing "ON" on the screen. Route the PLL control signal to one of the front panel outputs and choose the scale factor in V/Hz. I chose the same number as I measured for the DFD (including the SR560 gain) for ease of comparison on the spectrum analyzer.
I'm not sure what to believe. One would think the Zurich PLL is the most trustworthy, but a) I still am bothered by the time-domain behavior I see in the PLL control signal when I adjust the laser beat while watching it, and b) I've generated two nearly identical spectra that differ from it using completely different schemes from measurement to FFT.
All that said, I think the excess noise (and thanks to Dmass for saving me time by pointing this out) is just coming from the ThorLabs drivers, so this should be redone when we have our low-noise ones.
If the "locked indicator" light is not green on the Zurich (first tab, under "Reference", then what you get out is junk (e.g. you have unlocked the lock in, and i hasn't re-acquired yet) - you can do this by kicking it too hard with a frequency shift, which would be easy to do if you were slewing laser frequency, as the coefficients of the laser [Hz/mA] is so big. When the lock in loses the signal, you have to manually re-lock it (toggle off and on the button which has the mouseover text: "enable the fixed center frequency mode of the PLL"). You can get something which sort of looks like a PLL signal which has terrible noise and weird glitchy response when the lock in isn't locked in.
Your instinct to look for slewing at the PLL control point is correct, and a sign that the state of the PLL is healthy/unhealthy
Yes, I noticed this effect. I'm talking about immediately after acquiring---or re-aquiring---PLL lock. I did this several times at different beat frequencies to see what effect it had on the noise (the spectrum changed considerably, which is another bad sign).
On Monday, after I did some inventory of all the parts we have received from various companies, Dmass helped me mount the RIO lasers into their mounts so that I could get started with the optical setup. We cleaned the surfaces with methanol, applied a small layer of silver thermal compound, and then screwed them in.
I then borrowed the following to run the lasers:
After finding the right cables, I powered up the lasers and verified the P-I curve for each as listed on the spec sheets.
I then built a quick (temporary) optical beat setup, combining the two beams on an 1811. I had the temperatures (actually, thermistor resistances) set to what was listed as the testing set point on the datasheet, and as soon as I overlapped the beams and focused them onto the PD, there was already a strong ~50 MHz optical beat.
I have spent some time since then trying to lock various kinds of PLLs, both to interrogate the free-running frequency noise and to get used to controlling the lasers. Some things I've tried:
The first two were not helped by the fairly basic loop shaping afforded by attenuators and an SR560.
I think my next step will be to simply use the I-Q demodulation method (like I did to measure the no-FM Marconi noise in ATF:1877) to measure the frequency noise. I'll compare that to what I get with the Zurich PLL.
(I realized that we should probably use the CRYO elog rather than the SUS one, so I've reposted this here).
Today, we unpacked the IR Labs cryostat that will be the centerpiece of the Cryo SUS experiment.
Everything was more or less in order, except that the baseplate does not have any outward extensions with which to mount the cryostat to the table. Also, the holes for the screws holding the baseplate to the barrel are not countersunk. So, as of right now, the entire cryostat sits on these screws' caps, which is not ideal. We need to either a.) get a new baseplate made up with some wings on it and countersinking for the screws, or b.) work out another way to hold and mount the cryostat (for example, we might want some soft isolating material there anyway, though it will come at the expense of alignment drift).
I followed the instructions and removed the strange anodized heat shield bottom plate that comes with it during shipping, replacing it with the usual one and then resealing the chamber. As directed, I also pumped out the air again---the charcoal getter is not supposed to be exposed to atmosphere for long periods of time.
Photothermal (and other absorption effects) actually measured.
Cavity loss at 300K is 10ppm for west cavity. Fit looks very good, and jives with our understanding of the transfer functions. More to follow.
The 4395A network analyzer in the Cryo lab takes ages to save data to floppy. Someone had hooked up a Prologix GPIB-ethernet adapter to it, but it wasn't working on our network. I set it up as follows:
So, it acts like it has the static address 10.0.5.222 while on the cryo network -- but you can still take it to other networks and use it without monkeying with the settings.
Python scripts copied over from the 40m are installed on gaston, in the directory ~controls/netgpibdata. The AG4395A.py script was tested and works (much faster than the floppy drive).
While testing, Nic and I found and disabled a rogue DHCP server running on the framebuilder.
I think that if you see a signal in the demodulated PDH error point with the cavity unlocked, that this must be RF AM on the light.
What mechanism would produce this much AM? It can't be made through common path modulation of the carrier and sidebands. It must be an etalon formed somewhere between the EOM and the input mirror of the suspended cavity. Could be windows / viewports; this can be tested with the ND filter insertion technique we discussed on Wednesday.
The 4395 saved data in dBm/Hz while displaying Vrms/rtHz, so I had to figure out the conversion factor, and found the following useful table:
The Zurich Instruments HF2 has a very nice built in PLL feature. You give it a sine wave up to 50MHz and it will lock a PLL on it, and it will give the control signal at one of the outputs.
It has a nice PLL design interface (first attachment) (the design can be compared to the measured closed loop gain in the second attachment, this doesn’t include the marconi frequency modulation calibration). You tell it the bandwidth you want and it can internally set it’s PID to achieve that, and it will complain if you are asking it for too much. It claims to be able to get a 50kHz UGF, but I was only reliably able to get like 20kHz.
So right off the bat, this won’t be what we want if our goal is a very high BW PLL to suppress whatever nonlinear noise mechanism Rana and Dmass are worried about.
However, if 20kHz BW is enough, then this might be a pretty nice PLL to use. With a 1MHz modulation range, 20kHz PLL BW, and a 20MHz carrier, the noise is given in the last attachment. (below 1mHz/rtHz up to 1kHz, then starts to rise like f until it hits the PLL bandwidth, and it rolls off again.) The red trace is the spectrum analyzer noise.
data and scripts are all here.
PDH sensing noise:
New lower sensing noise with addition of transformers between LO and RF in both paths.
Cavity unlocked (detuned with temp) noise shows the scatter bump - unsure if this is interesting or should be totally obvious
The LB input noise gets to a minimum of 40 nV/rtHz if we turn the gain 100% up at high freq. This is 4x over what they claim
(shown it is 50 nV/rtHz)
Data is all on svn and will be put into noisebudget update